A 25.0-g block of ice at -15.00°C is dropped into a calorimeter (of negligible heat capacity) containing water at 15.00°C. When equilibrium is reached, the final temperature is 8.00°C. How much water did the calorimeter contain initially? The specific heat of ice is 2090 J/kg ∙ K, that of water is 4186 J/kg ∙ K, and the latent heat of fusion of water is 33.5 × 104 J/kg.
Answer #1

Specific heat of ice = C1 = 2090 J/(kg.K)

Specific heat of water = C2 = 4186 J/(kg.K)

Latent heat of fusion of water = L = 33.5 x 104 J/kg

Mas of the block of ice = m1 = 25 g = 0.025 kg

Mass of water in the calorimeter = m2

Initial temperature of the ice = T1 = -15 oC

Initial temperature of the water = T2 = 15 oC

Final equilibrium temperature = T3 = 8 oC

Melting point of ice = T4 = 0 oC

The heat energy gained by the ice is equal to the heat energy lost by the water.

m1C1(T4 - T1) + m1L + m1C2(T3 - T4) = m2C2(T2 - T3)

(0.025)(2090)(0 - (-15)) + (0.025)(33.5x104) + (0.025)(4186)(8 - 0) = m2(4186)(15 - 8)

m2 = 0.341 kg

Mass of water in the calorimeter initially = 0.341 kg